# The Cofinality of Generating Families

### v1

@@ Abstract
Let $\mathfrak{b}$ be the minimum size of an unbounded set in $\omega^{\omega}$ with the mod finite order. For a cardinal $\kappa$, the sampling number, $\mathop{sam}(\kappa)$, is the least number of countable subsets of $\kappa$ needed to have infinite intersection with every countably-infinite subset of $\kappa$. It is shown using the Tukey order on relations that (1) $\mathop{seq}(M)=\mathop{sam}(|M|)\cdot\mathfrak{b}$, unless $M$ is locally small (every point of $M$ has a neighborhood of size strictly less than $|M|$) in which case $\mathop{seq}(M)=\lim_{\mu<|M|}\mathop{sam}(\mu)\cdot\mathfrak{b}$ and (2) $k(M)$ is in the interval $kc(M)\cdot\mathfrak{b},\mathop{sam}(kc(M))\cdot\mathfrak{b}]$, where $kc(M)$ is the minimum number of compact sets that cover $M$. Shelah’s PCF theory is shown to provide tools to bound the sampling number, specifically, the covering number bounds from above, while mod finite scales give lower bounds.
@@ 1 Introduction
Van Douwen introduced the $k$-ness number in his survey article, [4], on cardinal characteristics of the continuum and small cardinals arising in topology. Observe that, for any separable metrizable space $M$, we have $k(M)\leq\mathop{seq}(M)\leq\mathfrak{c}$, the continuum ($\mathfrak{c}=|\mathbb{R}|=2^{\aleph_{0}}$). Van Douwen noted that $k(M)\leq\omega$ only when $M$ is locally compact, and he showed that when $M$ is not locally compact we have $\mathfrak{b}\leq k(M)$, where $\mathfrak{b}$ is the bounding number (minimum size of an unbounded set in $\omega^{\omega}$ with the mod finite order, $^{*}$). Van Douwen computed the $k$-ness number for a handful of spaces, notably the rationals, but left open (see [4, Question 8.11]) the computation of $k(M)$ for $M$ analytic or co-analytic. We conclude the paper with the solution to van Douwen’s question.
@@
To compute the sequentiality number, then, we need to understand the sampling number and its relation to the bounding number, $\mathfrak{b}$, and the continuum, $\mathfrak{c}$ (because separable metrizable spaces have cardinality no more than $\mathfrak{c}$). The basic properties and upper/lower bounds of the sampling number are studied in Section 3. This leads us to PCF theory (see Section 2.2 for background material on PCF theory). One of the fundamental objects of PCF theory is the covering number of a cardinal $\kappa$ which is the minimal size of a family of countable subsets of $\kappa$ such that every countable subset is contained in a member of the family. Always $\kappa\leq\mathop{cov}(\kappa)$, but $\mathop{cov}(\kappa)\geq\kappa^{+}$ for uncountable $\kappa$ with countable cofinality. It is a basic tenet of PCF theory that the covering number is ‘tame’ - $\mathop{cov}(\kappa)$ does not jump up much above $\kappa^{+}$. Indeed, at least assuming the Continuum Hypothesis, CH, it is known that large cardinal strength is required to get $\mathop{cov}(\kappa)>\kappa^{+}$. However, it is also known that, assuming the existence of a supercompact cardinal, it is consistent that $\mathop{cov}(\kappa)>\kappa^{+}$ for $\kappa=\aleph_{\omega}$. This is relevant because the sampling and covering numbers are clearly siblings. In particular, $\mathop{cov}(\kappa)$ is an upper bound of $\mathop{sam}(\kappa)$, and so upper bounds on the covering number give upper bounds on the sampling number. To show that the sampling number can also jump up, $\mathop{sam}(\kappa)>\kappa^{+}$, we show that the length of mod finite scales give lower bounds on $\mathop{sam}(\kappa)$, and show, modulo large cardinals, that it is consistent there are long scales. Section 3 concludes with a discussion of how simple forcing notions can change the covering and sampling numbers. This allows to construct, again modulo large cardinals, models where, for example, $\mathfrak{b}=\aleph_{1}$ (and so is a non-factor in the possible values of $\mathop{seq}$), $\mathop{sam}(\aleph_{\omega})>\aleph_{\omega+1}$ and $\mathfrak{c}$ is as large as we like. The concluding paragraph of Section 4.1 gives details on how our understanding of the sampling number informs our knowledge of the sequentiality number.
@@
Finally, in Section 4.2 we turn to the question of how many compact subsets suffice to generate the topology of a separable metrizable space. Let $kc(M)$, the compact covering number, be the minimum number of compact subsets needed to cover a separable metrizable $M$. The case of $\sigma$-compact spaces (i.e. $kc(M)\leq\omega$) was known to van Douwen, so we move on to $M$ with $kc(M)$ uncountable. For such $M$ it is no surprise that $kc(M)\leq k(M)$, but it was a surprise to the authors that $k(M)\leq\mathop{sam}(kc(M))\cdot\mathfrak{b}$, and so - as discussed above in the context of the sequentiality number - the compact covering and $k$-ness numbers are tightly coupled. But less so than the sequentiality number of a space and its cardinality. Recall for a fixed $\kappa$, $\mathop{seq}(M)$, for any $M$ with $|M|=\kappa$, is constrained to at most two values. However we show, modulo large cardinals, for every finite $K\geq 2$ it is consistent that, for a fixed $\kappa$, $k(M)$, for any $M$ with $kc(M)=\kappa$, takes on exactly $K$ values, and it is consistent that, for a fixed $\kappa$, there are infinitely many values of $k(M)$, for $M$ with $kc(M)=\kappa$. Section 4.2 concludes with the solution to van Douwen’s questions on the $k$-ness numbers of analytic and co-analytic spaces.
More on PCF and Open Questions
The connection between PCF theory and the sampling, sequentiality and $k$-ness numbers seems to the authors to be novel - and raises interesting questions - from at least three perspectives.
First, the sampling is clearly closely related to the covering number, but, formally at least, distinct from it. Hence the sampling number appears to be a new factor in PCF theory. However, the authors do not know how to distinguish the sampling and covering numbers: is $\mathop{sam}(\kappa)=\mathop{cov}(\kappa)$ for all $\kappa$? See Question C.
Second, the sequentiality and $k$-ness numbers depend on the sampling number of a cardinal $\kappa$ where $\kappa\leq\mathfrak{c}$. Historically, however, PCF theory has focused - motivated by the Singular Cardinals Hypothesis, for example - on singular cardinals in the presence of CH. So there is a mismatch between our needs and the literature. In particular, as mentioned above, assuming CH, it is known that large cardinal strength is required to get $\mathop{cov}(\kappa)>\kappa^{+}$. Can the Continuum Hypothesis be removed? See Question A.
Third, the lower bounds on the sampling number come from the additivity and existence of scales on products of cardinals with the mod finite order. PCF theory gets its very name from the cofinalities of products ordered modulo an ultrafilter, about which a great deal is known. Again there is a mismatch - mod finite versus modulo an ultrafilter. See Question B for details, and see also [9] for related problems. Because the covering number of the first fixed point of the $\aleph$-function can be arbitrarily large, lower bounds of the sampling number are of special interest in this case, see Question D.
@@ 2.1 Relations and Tukey Order
Let $(S_{n})_{n}$ be a sequence of sets. Let $\left(\prod_{n}[S_{n}]^{<\omega}\right)_{\infty}$ be all $(F_{n})_{n}$ in the product such that $F_{n}\neq\emptyset$ for infinitely many $n$. Define a general relation $i_{\infty}$ on sequences of sets by $(F_{n})_{n}i_{\infty}(G_{n})_{n}$ if and only if for infinitely many $n$ we have $F_{n}\cap G_{n}\neq\emptyset$. When all the $S_{n}$ are equal, say to $S$, then we write $\left([S]^{<\omega}\right)^{\omega}_{\infty}$ for $\left(\prod_{n}[S_{n}]^{<\omega}\right)_{\infty}$. For a cardinal $\kappa$, we now define the three relations $P(\kappa)$, $Q(\kappa)$, and $S(\kappa)$. Let $P(\kappa)=(\left([\kappa]^{<\omega}\right)^{\omega}_{\infty},i_{\infty})$. If $\operatorname{cof}(\kappa)\neq\omega$ then set $Q(\kappa)$ and $S(\kappa)$ to be the trivial relation. Otherwise, when $\operatorname{cof}(\kappa)=\omega$, fix a strictly increasing sequence of infinite cardinals, $(\kappa_{n})_{n}$ with limit $\kappa$, and set $Q(\kappa)=(\left(\prod_{n}[\kappa_{n}]^{<\omega}\right)_{\infty},i_{\infty})$ and $S(\kappa)=\sum_{n}P(\kappa_{n})\times(\kappa_{n},=)$. From [7, Lemma 2.14] we know $Q(\kappa)$ and $S(\kappa)$ are well-defined in the case when $\kappa$ has countable cofinality, because their Tukey type is independent of the choice of cofinal sequence, $(\kappa_{n})_{n}$. We know [7, Lemma 2.10] that $P(\omega)=_{T}(\omega^{\omega},\leq_{\infty})$, where $f\leq_{\infty}g$ if $\{n\in\omega:f(n)\leq g(n)\}$ is infinite, and hence $\operatorname{cof}(P(\omega))=\mathfrak{b}$. We end this subsection with the following simple lemma
@@ 2.2 PCF Theory
Relations of the form $([\mu]^{\delta},\subseteq)$, and their cofinalities, where $\mu$ is singular and $\delta\leq\mu$, are central objects of PCF theory. For the purposes of this paper, the case $\delta=\omega$ is the most important, and, simplifying terminology and notation, we define the covering number, $\mathop{cov}(\kappa)$, of a cardinal $\kappa$, to be $\operatorname{cof}([\kappa]^{\omega},\subseteq)$.
@@
(Gitik, [12]) Assuming the existence of certain large cardinals, $\mathop{cov}(\kappa)$ can be arbitrarily large, where $\kappa$ is the first fixed point of the $\aleph$-function.
@@
Mod Finite Scales
Fix a strictly increasing sequence $\left<\kappa_{n}:n\in\omega\right$> of regular cardinals, say with limit $\kappa$. Let $^{*}$ be the mod finite order on $\prod_{n}\kappa_{n}$, so $f<^{*}g$ if and only if for all but finitely many $n$ we have $f(n)<g(n)$. A (mod finite) scale in $\prod_{n\in\omega}\kappa_{n}$ is a sequence $\vec{f}=\left<f_{\alpha}:\alpha<\lambda\right$> of functions in the product which is increasing mod finite ($\alpha<\beta<\lambda$ implies $f_{\alpha}<^{*}f_{\beta}$) and cofinal mod finite. The cardinal $\lambda$ is called the length of the scale. In Tukey terms, there is a scale of length $\lambda$ in $\prod_{n}\kappa_{n}$ if and only if $\left(\prod_{n}\kappa_{n},<^{*}\right)=_{T}\lambda$.
Shelah (see Jech, [14] for a proof) showed that if $\kappa$ is an uncountable cardinal of countable cofinality then there exists a strictly increasing sequence $\left<\kappa_{n}:n\in\omega\right$> of regular cardinals, with limit $\kappa$, such that there is a scale on $\prod_{n\in\omega}\kappa_{n}$ of length $\kappa^{+}$.
3 The Sampling Number
3.1 Definition and Basic Properties
Let $\operatorname{ii}$ be the general relation ($\operatorname{ii}$ for “infinite intersection”) defined as follows: $A\operatorname{ii}B$ if and only if $A\cap B$ is infinite. Define the sampling number of a cardinal $\kappa$ to be $\mathop{sam}(\kappa)=\operatorname{cof}([\kappa]^{\omega},\operatorname{ii})$. That is, $\mathop{sam}(\kappa)$ is the least number of countable subsets of $\kappa$ needed to have infinite intersection with every countably-infinite subset of $\kappa$. For $\kappa\geq\lambda$ the map $\phi_{+}:[\kappa]^{\omega}\to[\lambda]^{\omega}$ given by restriction, $\phi_{+}(x)=x\cap\lambda$, carries cofinal sets to cofinal sets, so $\mathop{sam}$ is monotonic.
Lemma 3.1.
If $\kappa\geq\lambda$, then $\mathop{sam}(\kappa)\geq\mathop{sam}(\lambda)$.
Lemma 3.2.
Let $\kappa$ be an uncountable cardinal. Then
  1. (i)
    $\kappa\leq\mathop{sam}(\kappa)\leq\kappa^{\omega}$, and $\mathop{sam}(\kappa^{\omega})=\kappa^{\omega}$,
  2. (ii)
    if $\operatorname{cof}(\kappa)=\omega$ then $\mathop{sam}(\kappa)>\kappa$,
  3. (iii)
    if $\kappa=\mu^{+}$ then $\mathop{sam}(\kappa)=\mathop{sam}(\mu)\cdot\kappa$, and
  4. (iv)
    if $\kappa$ is a limit and $\operatorname{cof}(\kappa)>\omega$ then $\mathop{sam}(\kappa)=\lim\{\mathop{sam}(\mu):\mu<\kappa\}$.
Proof.
(i) Let $\mathcal{C}\subseteq[\kappa]^{\omega}$ with $|\mathcal{C}|<\kappa$. Then $S=\kappa\setminus\bigcup\mathcal{C}$ has size $\kappa$. Hence any $B$ in $S]^{\omega}$ is disjoint from every $A$ in $\mathcal{C}$, and so the latter is not cofinal in $([\kappa]^{\omega},\operatorname{ii})$. So $\kappa\leq\mathop{sam}(\kappa)$. As $|[\kappa]^{\omega}|=\kappa^{\omega}$, clearly, $\mathop{sam}(\kappa)\leq\kappa^{\omega}$. Now $\mathop{sam}(\kappa^{\omega})=\kappa^{\omega}$ follows.
(ii) Fix a sequence of cardinals, $(\kappa_{n})_{n}$, strictly increasing to $\kappa$. Let $\mathcal{C}\subseteq[\kappa]^{\omega}$ with $|\mathcal{C}|\leq\kappa$, say, $\mathcal{C}=\{A_{\alpha}:\alpha<\kappa\}$. For each $n$, let $\mathcal{C}_{n}=\{A_{\alpha}:\alpha\leq\kappa_{n}\}$, and set $C_{n}=\bigcup\mathcal{C}_{n}$. Note that the $C_{n}$’s are increasing. Since, for each $n$, $|C_{n}|\leq\kappa_{n}<\kappa$, we can pick $\alpha_{n}\in\kappa\setminus(C_{n}\cup\left\{\alpha_{m}:m<n\right\})$. Let $B=\{\alpha_{n}:n\in\omega\}$, so $B$ is a countably infinite subset of $\kappa$. Take any $A_{\alpha}$ in $\mathcal{C}$. Then for some $m$, $A_{\alpha}$ is in $\mathcal{C}_{m}$, and $A_{\alpha}\subseteq C_{m}$. For any $n\geq m$, $\alpha_{n}\in\kappa\setminus C_{n}\subseteq\kappa\setminus C_{m}$, so $\alpha_{n}$ is not in $A_{\alpha}$. Hence $|B\cap A_{\alpha}|\leq m$, and, in particular, is finite. It follows that $\mathcal{C}$ is not cofinal in $([\kappa]^{\omega},\operatorname{ii})$.
(iii) Recalling that $\kappa=\mu^{+}$, let $\lambda=\mathop{sam}(\mu)$, and let $\{A_{\alpha}:\alpha<\lambda\}$ enumerate a cofinal set of $([\mu]^{\omega},\operatorname{ii})$. For each $\delta\in(\mu,\mu^{+})$, let $f_{\delta}$ be a bijection from $\mu$ onto $\delta$. Define $\mathcal{C}^{+}=\left\{f_{\delta}[A_{\alpha}]:\alpha<\lambda\text{ and }\delta\in(\mu,\mu^{+})\right\}$. Then $\mathcal{C}^{+}$ is cofinal in $([\mu^{+}]^{\omega},\operatorname{ii})$. To see this, take any $S\in[\mu^{+}]^{\omega}$, and pick $\delta<\mu^{+}$ so that $S\subseteq\delta$. Fix $\alpha<\lambda$ so that $f_{\delta}^{-1}[S]\cap A_{\alpha}$ is infinite. Then $f_{\delta}[f_{\delta}^{-1}[S]\cap A_{\alpha}]=S\cap f_{\delta}[A_{\alpha}]$ is infinite. Since $\mathop{sam}(\mu)$ and $\kappa$ are both $\leq\mathop{sam}(\kappa)$, equality follows.
(iv) By Lemma 3.1, we only need to show that $\mathop{sam}(\kappa)\leq\lim\{\mathop{sam}(\mu):\mu<\kappa\}$. For each cardinal $\mu<\kappa$, fix $\mathcal{C}_{\mu}$ cofinal in $([\mu]^{\omega},\operatorname{ii})$ of minimum size, $\mathop{sam}(\mu)$. Then $\mathcal{C}=\bigcup\{\mathcal{C}_{\mu}:\mu<\kappa\}$ has size $\lim\{\mathop{sam}(\mu):\mu<\kappa\}$, and is cofinal in $([\kappa]^{\omega},\operatorname{ii})$, because any countable subset of $\kappa$ is contained in some $\mu<\kappa$. ∎
Observe that $\mathop{sam}(\aleph_{1})=\aleph_{1}$ because $\mathcal{C}=\{[0,\alpha]:\alpha\in\omega_{1}\}$ is cofinal in $([\omega_{1}]^{\omega},\operatorname{ii})$. By induction, applying claims (i) and (iii) of the preceding lemma, we compute the sampling number of all cardinals strictly less than $\aleph_{\omega}$.
Lemma 3.3.
For every $n\in\mathbb{N}$ we have $\mathop{sam}(\aleph_{n})=\aleph_{n}$.
3.2 Upper Bounds - The Covering Number
Since cofinal sets for $([\kappa]^{\omega},\subseteq)$ are cofinal in $([\kappa]^{\omega},\operatorname{ii})$ we immediately see that the covering number is an upper bound of the sampling number.
Lemma 3.4.
Let $\kappa$ be an uncountable cardinal. Then $\mathop{sam}(\kappa)\leq\mathop{cov}(\kappa)$.
From the known upper bounds on the covering number we deduce upper bounds for the sampling number. Specifically, from Proposition 2.2 we see:
Corollary 3.5.
If $\kappa=\aleph_{\alpha}$ is an uncountable cardinal below the least fixed point of the $\aleph$-function, then $\mathop{sam}(\kappa)<\aleph_{|\alpha|^{+4}}$.
While from Theorem 2.5 we have:
Corollary 3.6.
$\mathop{sam}(\kappa)=\begin{cases}\kappa^{+}&\text{ if }\operatorname{cof}(\kappa)=\omega\\ \kappa&\text{ otherwise.}\end{cases}$
@@ 3.2 Upper Bounds - The Covering Number
In the context of computing the sequentiality and $k$-ness numbers, our interest is in $\mathop{sam}(\kappa)$ where $\kappa$ (the size of a separable metrizable space) is no more than $\mathfrak{c}$. Unfortunately, then, the preceding theorem does not apply.
@@
    If there is no $\lambda$ with Mitchell order $\lambda^{++}$, then is $\mathop{sam}(\aleph_{\omega})=\aleph_{\omega+1}$? Is $\mathop{cov}(\aleph_{\omega})=\aleph_{\omega+1}$?
@@
    Next we have a version of the previous item, but for other cardinals than $\aleph_{\omega}$ specifically: suppose that there is no $\lambda$ with Mitchell order $\lambda^{++}$, and that $\mathfrak{c}>>\aleph_{\omega}$. Is it true that for every $\omega<\kappa<\mathfrak{c}$ of countable cofinality, $\mathop{sam}(\kappa)=\kappa^{+}$? Is $\mathop{cov}(\kappa)=\kappa^{+}$?

### v2

@@ Abstract
Let $\mathfrak{b}$ be the minimum size of an unbounded set in $\omega^{\omega}$ with the mod finite order. For a cardinal $\kappa$, the covering number, $\mathop{cov}(\kappa)$, is the minimum size of a family of countable subsets of $\kappa$ so that every countable subset of $\kappa$ is contained in an element of the family. It is shown using the Tukey order on relations that (1) $\mathop{seq}(M)=\mathop{cov}(|M|)\cdot\mathfrak{b}$, unless $M$ is locally small (every point of $M$ has a neighborhood of size strictly less than $|M|$) in which case $\mathop{seq}(M)=\lim_{\mu<|M|}\mathop{cov}(\mu)\cdot\mathfrak{b}$ and (2) $k(M)$ is in the interval $kc(M)\cdot\mathfrak{b},\mathop{cov}(kc(M))\cdot\mathfrak{b}]$, where $kc(M)$ is the minimum number of compact sets that cover $M$.
@@ 1 Introduction
Van Douwen introduced the $k$-ness number in his survey article, [5], on cardinal characteristics of the continuum and small cardinals arising in topology. Observe that, for any separable metrizable space $M$, we have $k(M)\leq\mathop{seq}(M)\leq\mathfrak{c}$, the continuum ($\mathfrak{c}=|\mathbb{R}|=2^{\aleph_{0}}$). Van Douwen noted that $k(M)\leq\omega$ only when $M$ is locally compact, and he showed that when $M$ is not locally compact we have $\mathfrak{b}\leq k(M)$, where $\mathfrak{b}$ is the bounding number (minimum size of an unbounded set in $\omega^{\omega}$ with the mod finite order, $^{*}$). Van Douwen computed the $k$-ness number for a handful of spaces, notably the rationals, but left open (see [5, Question 8.11]) the computation of $k(M)$ for $M$ analytic or co-analytic. We conclude the paper with the solution to van Douwen’s question.
@@
To compute the sequentiality number, then, we need to understand the sampling number and its relation to the bounding number, $\mathfrak{b}$, and the continuum, $\mathfrak{c}$ (because separable metrizable spaces have cardinality no more than $\mathfrak{c}$). One of the fundamental objects of PCF theory is the covering number of a cardinal $\kappa$ which is the minimal size of a family of countable subsets of $\kappa$ such that every countable subset is contained in a member of the family. Clearly the sampling and covering numbers are closely related. In fact, see Theorem 2.16, despite the formal difference between them, they coincide: for every $\kappa$ we have $\mathop{sam}(\kappa)=\mathop{cov}(\kappa)$. From PCF theory much is known about the covering number and we recapitulate this next. However our results above on the sequentiality number, as well as those below on the $k$-ness number, are most naturally established for the sampling number. Since the sampling and covering numbers are equal, we can substitute one for the other at our convenience. For consistency we have stated our results below using the covering number, but in proofs often use the sampling number instead without explicitly (and repeatedly) stating the equality. The concluding paragraph of Section 3.1 gives details on how our understanding of the covering number informs our knowledge of the sequentiality number.
@@
It is known that $\kappa\leq\mathop{cov}(\kappa)$, but $\mathop{cov}(\kappa)\geq\kappa^{+}$ for uncountable $\kappa$ with countable cofinality. Further, it is a basic tenet of PCF theory that the covering number is ‘tame’ - $\mathop{cov}(\kappa)$ does not jump up much above $\kappa^{+}$. Indeed, at least assuming the Continuum Hypothesis, CH, large cardinal strength is required to get $\mathop{cov}(\kappa)>\kappa^{+}$. However, it is also known that, assuming the existence of a supercompact cardinal, it is consistent that CH holds and $\mathop{cov}(\kappa)>\kappa^{+}$ for $\kappa=\aleph_{\omega}$. Applying random real forcing to the latter model, allows us to construct, again modulo large cardinals, models where, for example, $\mathfrak{b}=\aleph_{1}$ (and so is a non-factor in the possible values of $\mathop{seq}$), $\mathop{cov}(\aleph_{\omega})>\aleph_{\omega+1}$ and $\mathfrak{c}$ is as large as we like. Section 2.2 provides details on the above, and concludes with a discussion of the cofinality of products of countable sets of cardinals with the mod finite order, and in particular the existence of mod finite scales. This enables the construction of models, modulo large cardinals, in which we can control the covering numbers of families of cardinals.
@@
Finally, in Section 3.2 we turn to the question of how many compact subsets suffice to generate the topology of a separable metrizable space. Let $kc(M)$, the compact covering number, be the minimum number of compact subsets needed to cover a separable metrizable $M$. The case of $\sigma$-compact spaces (i.e. $kc(M)\leq\omega$) was known to van Douwen, so we move on to $M$ with $kc(M)$ uncountable. For such $M$ it is no surprise that $kc(M)\leq k(M)$, but it was a surprise to the authors that $k(M)\leq\mathop{cov}(kc(M))\cdot\mathfrak{b}$, and so - as discussed above in the context of the sequentiality number - the compact covering and $k$-ness numbers are tightly coupled. But less so than the sequentiality number of a space and its cardinality. Recall for a fixed $\kappa$, $\mathop{seq}(M)$, for any $M$ with $|M|=\kappa$, is constrained to at most two values. However we show, modulo large cardinals, for every finite $K\geq 2$ it is consistent that, for a fixed $\kappa$, $k(M)$, for any $M$ with $kc(M)=\kappa$, takes on exactly $K$ values, and it is consistent that, for a fixed $\kappa$, there are infinitely many values of $k(M)$, for $M$ with $kc(M)=\kappa$. Section 3.2 concludes with the solution to van Douwen’s questions on the $k$-ness numbers of analytic and co-analytic spaces.
@@ 2.1 Relations and Tukey Order
Let $(S_{n})_{n}$ be a sequence of sets. Let $\left(\prod_{n}[S_{n}]^{<\omega}\right)_{\infty}$ be all $(F_{n})_{n}$ in the product such that $F_{n}\neq\emptyset$ for infinitely many $n$. Define a general relation $i_{\infty}$ on sequences of sets by $(F_{n})_{n}i_{\infty}(G_{n})_{n}$ if and only if for infinitely many $n$ we have $F_{n}\cap G_{n}\neq\emptyset$. When all the $S_{n}$ are equal, say to $S$, then we write $\left([S]^{<\omega}\right)^{\omega}_{\infty}$ for $\left(\prod_{n}[S_{n}]^{<\omega}\right)_{\infty}$. For a cardinal $\kappa$, we now define the three relations $P(\kappa)$, $Q(\kappa)$, and $S(\kappa)$. Let $P(\kappa)=(\left([\kappa]^{<\omega}\right)^{\omega}_{\infty},i_{\infty})$. If $\operatorname{cof}(\kappa)\neq\omega$ then set $Q(\kappa)$ and $S(\kappa)$ to be the trivial relation. Otherwise, when $\operatorname{cof}(\kappa)=\omega$, fix a strictly increasing sequence of infinite cardinals, $(\kappa_{n})_{n}$ with limit $\kappa$, and set $Q(\kappa)=(\left(\prod_{n}[\kappa_{n}]^{<\omega}\right)_{\infty},i_{\infty})$ and $S(\kappa)=\sum_{n}P(\kappa_{n})\times(\kappa_{n},=)$. From [8, Lemma 2.14] we know $Q(\kappa)$ and $S(\kappa)$ are well-defined in the case when $\kappa$ has countable cofinality, because their Tukey type is independent of the choice of cofinal sequence, $(\kappa_{n})_{n}$. We know [8, Lemma 2.10] that $P(\omega)=_{T}(\omega^{\omega},\leq_{\infty})$, where $f\leq_{\infty}g$ if $\{n\in\omega:f(n)\leq g(n)\}$ is infinite, and hence $\operatorname{cof}(P(\omega))=\mathfrak{b}$. We end this subsection with the following simple lemma
@@ 2.2 PCF Theory
Relations of the form $([\mu]^{\delta},\subseteq)$, and their cofinalities, where $\mu$ is singular and $\delta\leq\mu$, are central objects of PCF theory. For the purposes of this paper, the case $\delta=\omega$ is the most important, and, simplifying terminology and notation, we define the covering number, $\mathop{cov}(\kappa)$, of a cardinal $\kappa$, to be $\operatorname{cof}([\kappa]^{\omega},\subseteq)$. Basic facts about the covering number are recorded below, see [18].
Lemma 2.2.
If $\kappa\geq\lambda$, then $\mathop{cov}(\kappa)\geq\mathop{cov}(\lambda)$.
Lemma 2.3.
Let $\kappa$ be an uncountable cardinal. Then
  1. (i)
    $\kappa\leq\mathop{cov}(\kappa)\leq\kappa^{\omega}$, and $\mathop{cov}(\kappa^{\omega})=\kappa^{\omega}$,
  2. (ii)
    if $\operatorname{cof}(\kappa)=\omega$ then $\mathop{cov}(\kappa)>\kappa$,
  3. (iii)
    if $\kappa=\mu^{+}$ then $\mathop{cov}(\kappa)=\mathop{cov}(\mu)\cdot\kappa$, and
  4. (iv)
    if $\kappa$ is a limit and $\operatorname{cof}(\kappa)>\omega$ then $\mathop{cov}(\kappa)=\lim\{\mathop{cov}(\mu):\mu<\kappa\}$.
Lemma 2.4.
For every $n\in\mathbb{N}$ we have $\mathop{cov}(\aleph_{n})=\aleph_{n}$.
@@
(Gitik, [13]) Assuming the existence of certain large cardinals, $\mathop{cov}(\kappa)$ can be arbitrarily large, where $\kappa$ is the first fixed point of the $\aleph$-function, and CH holds.
@@
In the context of computing the sequentiality and $k$-ness numbers, our interest is in $\mathop{cov}(\kappa)$ where $\kappa$ (the size of a separable metrizable space) is no more than $\mathfrak{c}$. Unfortunately, then, the preceding theorem does not apply.
@@
    If there is no $\lambda$ with Mitchell order $\lambda^{++}$, then is $\mathop{cov}(\aleph_{\omega})=\aleph_{\omega+1}$?
@@
    Next we have a version of the previous item, but for other cardinals than $\aleph_{\omega}$ specifically: suppose that there is no $\lambda$ with Mitchell order $\lambda^{++}$, and that $\mathfrak{c}>>\aleph_{\omega}$. Is it true that for every $\omega<\kappa<\mathfrak{c}$ of countable cofinality, $\mathop{cov}(\kappa)=\kappa^{+}$?